OpenStudy (anonymous):

What is the derivative of y = 1/ x+radx

hartnn (hartnn):

its this \(\frac{1}{x+\sqrt x}\) right ?

hartnn (hartnn):

if yes then , use quotient rule.....

OpenStudy (anonymous):

y' = -1/(x+radx)^2 * (1+1/2radx)

hartnn (hartnn):

yes, even thats correct..

OpenStudy (anonymous):

\[y'=(\frac{ 1 }{ (x+\sqrt{x})^2 })(1+\frac{ 1 }{ 2\sqrt{x} })\] Would yo need to simplify any further?

OpenStudy (anonymous):

Or, could you?

hartnn (hartnn):

no, simplifying further would not help... all you can do is \(y'=(\frac{ 1 }{ (x+\sqrt{x})^2 })(\frac{ 1+2\sqrt x }{ 2\sqrt{x} })\) which is not much simplified

OpenStudy (anonymous):

Alright, thank you very much!

hartnn (hartnn):

welcome :)